# AuON: A Linear-time Alternative to Semi-Orthogonal Momentum Updates

### v2

@@ 4 Experiments
4.1 Language Modeling
We evaluate our approach using a 4X L4 GPU on the SmolLM-Corpus datasetBen Allal et al. (2024) , consisting of 500k tokens. The underlying model is a nanoGPTKarpathy (2022) with FlashAttention-2Dao (2023) rotary position embeddings (RoPE)Su et al. (2023), RMSNorm Huang et al. (2019), and SwiGLU activations Shazeer (2020). For the Small configuration, we use a hidden size of 512, 6 layers, 8 attention heads, and a feed-forward dimension of 1536. Training is conducted for 6000 steps with a global batch size of 128. We compare AuON, AdamW Loshchilov and Hutter (2019), Hybrid-AuON, and MuON under similar training conditions, with learning rates tuned separately: $\eta_{\text{adamw}}=0.003$, $\eta_{\text{auon}}=0.24$. and $\eta_{\text{muon}}=0.01$. Rosić and Claude (2025)
@@ 4.2 Combined Analysis: MNIST and CIFAR-10 CNN
We conduct a combined evaluation of AuON, SGD, and AdamW on two vision benchmarks: a simple multilayer perceptron (MLP) on MNIST and a convolutional neural network (CNN) on CIFAR-10. These experiments, adapted from Wei (2025) Wei (2025), are designed to assess both convergence behavior and runtime efficiency in progressively more challenging tasks.
@@ 5 Conclusion
We introduced AuON, a linear-time optimizer that enforces unit spectral norm constraints to stabilize training without the computational overhead of full semi-orthogonalization. Across tasks of varying complexity—from nanoGPT on SmolLM-Corpus to MNIST and CIFAR-10—AuON consistently improved convergence and accuracy over SGD, while matching or surpassing AdamW at nearly identical runtime cost. The hybrid variant further narrows the gap to MuON, offering a lighter-weight alternative with strong empirical performance.
Despite these gains, AuON and Hybrid-AuON may exhibit instability on very large-scale transformer models, such as exploding attention logits, as noted in prior work
Team et al. (2025). Due to limited GPU resources, we leave large-scale evaluations to future work, where techniques like QK-clipping may help mitigate these issues.
Overall, our findings indicate that AuON is a practical and efficient alternative to standard optimizers and a lightweight counterpart to MuON, capable of delivering robust optimization performance without sacrificing efficiency. Future work will focus on scaling AuON to larger architectures and long-horizon training runs, such as NanoGPT speedruns on H100 GPUs, to further establish its robustness in production-scale training.
@@ Appendix A.1  Convergence Guarantee for AuON
where NN is the update dimension.
@@ Appendix A.2  Experimental Validation of Assumptions
Table 5 reports representative values at different training steps. We observe that the training loss decreases monotonically, $\hat{\kappa}$ remains strongly positive (median values between $8.05$ and $13.66$, with 10th-percentiles always above 77), and $\hat{\sigma}^{2}$ is tightly bounded near $1.0$. This matches the theoretical spectral trust-region guarantee of AuON: updates remain well-aligned while their variance is constrained to a stable unit scale.
@@ 28 U = update / (rms + 1e-8)
33class Muon(torch.optim.Optimizer):

### v3

@@ 3.4 Spectral Trust-Region Guarantee and Reduction of Correlation Energy in AuON
3.5 Emergent “Emergency Brake” Behavior in AuON via Hyperbolic-Cosine Normalization
A critical failure mode in large-scale Transformer training is the phenomenon of “exploding at tention logits.” As weight norms growunbounded, thequery-keydotproductsdiverge, causing Softmax saturation. This results in gradient matrices that are geometrically sparse and heavy tailed (high kurtosis).
AuON introduces a hyperbolic-cosine–based normalization that acts as an automatic “emergency brake” when gradients become unstable, for example under exploding attention heads. Unlike Muon, which normalizes to unit norm regardless of the gradient distribution, AuON computes a scalar factor
3.5.1 Cosh-Based Normalization as an Instability Detector: The Hyperbolic Sensor
We demonstrate that AuON possesses an inherent self-regulating mechanism—the Emergency Brake—which is an emergent property of the hyperbolic cosine scaling. Unlike semi-orthogonal projections that enforce unit energy updates regardless of landscape curvature, AuON utilizes the $\cosh(\cdot)$ function as a non-linear stability sensor.
The core stability mechanism lies in the construction of the normalization scalar $r$. Recall the temperature-scaled definition:
$r=\sqrt{\frac{1}{N}\sum_{i,j}\cosh^{2}\left(\gamma\cdot\tilde{G}_{ij}\right)}\quad\text{where }\gamma=\sqrt{N}$ (6)
The function $f(x)=\cosh(x)$ exhibits two distinct asymptotic behaviors depending on the magnitude of the input gradient:
  1. 1.
    The Safe Zone (Quadratic Regime): For well-behaved gradients where values are near zero, $\cosh(x)\approx 1+x^{2}/2$. In this regime, $r$ behaves similarly to a standard $L_{2}$ norm, facilitating efficient optimization.
  2. 2.
    The Danger Zone (Exponential Regime): For large outliers (spikes) caused by exploding logits, $\cosh(x)\approx\frac{1}{2}e^{|x|}$.
Consider a scenario where the gradient matrix develops an instability spike of magnitude $\xi\gg 1$. The normalization factor $r$ grows exponentially:
$r\propto e^{|\xi|}$ (7)
[…]
$U||_{2}\propto\frac{\xi}{e^{|\xi|}}\xrightarrow{\xi\to\infty}0$ (8)
This proves that AuON automatically throttles the update size towards zero as instability increases, effectively freezing the weights of the unstable layer to allow for energy dissipation. In effect, cosh acts as a nonlinear sensor that is nearly quadratic in the “safe” regime but becomes extremely sensitive to rare, large-magnitude components.
3.5.2 Sequence of Brake Activation Under Exploding Attention
When an attention head becomes unstable, attention logits can grow excessively large, driving the softmax into saturation and producing a gradient matrix $\tilde{G}$ that is sparse and spiky: most entries remain near zero, while a small subset corresponding to unstable tokens attain very high magnitudes (e.g., on the order of $10$ after pre-normalization).
In conventional $\ell_{2}$-based normalization, a single spike of magnitude $10$ contributes only $10^{2}=100$ to the norm sum over millions of parameters—negligible in the global scale and therefore ineffective for damping such instabilities.
Under AuON, the same spike passes through $\cosh$: for $\tilde{G}_{ij}=10$, $\cosh(10)$ is on the order of $10^{4}$, and AuON further squares this, yielding $\cosh^{2}(10)\approx 10^{8}$. Even after averaging over all $N$ parameters, this single outlier term dominates the computation of $r$, causing the normalization factor to grow dramatically relative to its normal value.
3.5.3 Global Suppression as the “Emergency Brake”
The final AuON update is defined as
$U=\frac{\tilde{G}}{r}.$ (9)
Once a spike drives $r$ to a large value, the entire update matrix $U$ is proportionally downscaled. The offending spiky entries are strongly suppressed, and the otherwise normal entries are simultaneously reduced toward zero, effectively freezing that layer’s update for the current step. This global suppression acts as an automatic emergency brake: it activates only when the gradient field contains extreme outliers yet leaves normal training dynamics largely unaffected when all entries of $\tilde{G}$ lie within the quasi-quadratic region of $\cosh$.
3.5.4 Empirical Verification at Production Scale (d=4096)
To stress-test the “Emergency Brake” mechanism at a realistic scale, we constructed a Transformer layer with model dimension d = 4096 and an MLP expansion factor of 4(Dmlp = 4096), totaling approximately 100.6 Million million parameters. Crucially, all normalization layers (LayerNorm/RM SNorm) were removed and the learning rate was kept at 0.24 . This architecture creates a high-pressure optimization environment where weight accumulation leads directly to attention logit explosion. We compared the Newton-Schulz semi-orthogonal update (Muon) against AuON (with temperature scaling )over 1000 steps.
Figure 3: The Emergency Brake in Action. The Muon optimizer (Red) exhibits step-function explosions, driving logits to hard saturation (greater than 30). In contrast, AuON (Blue) and AuON+5(Ns=5)(Green) demonstrat the theoretically predicted logarithmic braking curve, capping logit growth at approximately 12 even after 1,000 steps of unnormalized accumulation.
As illustrated in Figure 3, the divergence in behavior is stark:
  • Muon (No Brake): The semi-orthogonal update ignores the growing gradient magnitude, treating the exploding logits as a strong signal. This leads to immediate saturation and chaotic quantization artifacts.
  • AuON and Hybrid-AuON(Ns=5) (Brake Engaged): The blue and green trajectorys confirms the exponential suppression derived in Eq. (9). As logits attempt to grow, the gradient kurtosis rises, triggering the $\cosh$ term to inflate the denominator $r$. The result is a smooth, controlled saturation curve rather than a catastrophic explosion.
3.5.5 Long-Horizon Stability: The Dissipative Property
While short-term stress tests reveal the immediate reaction to shock, long-horizon training reveals the fundamental thermodynamic properties of the optimizer. We extended the unnormalized Transformer stress test ($d=1024$) to 20,000 steps to observe the asymptotic behavior of the logit magnitudes.
Figure 4: Thermodynamics of Optimization (20k Steps). The Muon optimizer (Red) exhibits high-variance chaotic behavior with a persistent upward drift, indicative of unregulated energy injection. In stark contrast, AuON (Blue) demonstrates a dissipative trajectory. After an initial rise due to the lack of normalization, the self-regulating cosh mechanism forces a inflection point at step approximately 2,500, after which the logit magnitudes steadily decay. This proves that AuON actively removes instability energy from the system over time.
3.5.6 Analysis of Asymptotic Behavior
The trajectory in Figure 4 highlights a critical distinction in stability mechanics:
  1. 1.
    Chaotic Energy Injection (Muon): The high variance (jaggedness) of the red curve indicates that Muon is continuously oscillating between saturation states. The spectral constraint (update norm equals 1) forces the optimizer to maintain high kinetic energy even when the loss landscape suggests caution, preventing convergence to a stable basin.
  2. 2.
    Controlled Energy Dissipation (AuON): The smooth, downward trajectory of the blue curve confirms the theoretical prediction of Section 3.3. As logits grow, the gradient kurtosis increases. The temperature-scaled cosh denominator ($r$) responds exponentially, driving the effective update norm $\|U\|_{2}\ll 1$. This turns the optimizer into an energy sink, draining the excess weight magnitude accumulated during the early training phase and naturally guiding the system back toward a stable regime ($\approx 9.5$) despite the absence of explicit normalization layers.
3.6 Empirical Validation: The Geometry of Harmful Directions
To validate the hypothesis that harmful optimization directions are characterized by spatial spikiness (high kurtosis), we analyzed both real-time training dynamics and static landscape geometry. The results provide strong empirical confirmation that spikiness serves as a reliable geometric signature of instability.
3.6.1 Dynamic Instability Analysis (The “Smoking Gun”)
Our primary evidence is derived from monitoring the Fisher kurtosis of gradient updates during the training of a Transformer model. As illustrated in Figure 5, we observe a strong temporal correlation between geometric spikiness and loss instability.
Figure 5: Training Monitor: Correlation between loss instability (red) and gradient kurtosis (blue). Figure 1: Temporal correlation of instability. The training trajectory reveals that significant spikes in training loss (instability events, e.g., at steps 24, 40, and 46) are inextricably linked to simultaneous or preceding spikes in gradient kurtosis. This establishes high kurtosis as a leading indicator of divergence in real-world optimization.
The synchronization between loss spikes (red line) and gradient kurtosis (blue dashed line) confirms that when the optimizer naturally encounters “bad directions,” those directions manifest as heavy-tailed, spiky distributions. This validates the premise that the onset of exploding attention logits creates an impulsive noise signature that can be detected and suppressed via kurtosis-sensitive normalization.
3.6.2 Geometric Risk Analysis (The Iso-Norm Probe)
To disentangle the effects of vector geometry from vector magnitude, we conducted a landscape probe using iso-norm perturbations. We injected random noise vectors—normalized to identical $\ell_{2}$ norms but differing in spatial distribution—into a converged model. Figure 6 presents the comparative impact on validation loss.
Risk Distribution.
The analysis reveals a nuanced but critical finding regarding optimization risk:
  • Average vs. Variance: While the average harm of random spiky directions is comparable to that of flat directions, the variance is significantly higher. Flat directions represent “safe bets” with predictable effects on loss.
Figure 6: Landscape Probe: Loss increase vs. direction spikiness for iso-norm vectors. Figure 2: The geometry of risk. Gaussian (flat) perturbations (blue circles) cluster tightly with low impact, while spiky perturbations (red triangles) exhibit higher variance. Crucially, the most damaging directions (highest $y$-axis values) are exclusively spiky.
Maximum Risk Confirmation.
The single most damaging direction identified in the experiment (causing a loss increase of +0.0025+0.0025) corresponded to a high-kurtosis vector. This confirms that while not all spiky directions are harmful, the most dangerous outliers in the loss landscape—those corresponding to the sharpest curvature—are statistically aligned with spiky vectors.
3.6.3 Summary of Findings
Combining the dynamic evidence (Test B) with the static probe (Test A), we conclude that spikiness is the geometric signature of harmful optimization directions. Flat (Gaussian) updates represent the dense, informative signal required for learning, while high-kurtosis updates represent high-risk instability.
The AuON optimizer’s “emergency brake” mechanism effectively mitigates this risk by selectively suppressing the high-variance, spiky components identified in Figure 6, thereby preventing the divergence events visualized in Figure 5.
@@ 4 Experiments
4.1 Small Language Modeling
We evaluate our approach using a 4X L4 GPU on the SmolLM-Corpus datasetbenallal2024smollmcorpus , consisting of 500k tokens. The underlying model is a nanoGPTKarpathy2022 with FlashAttention-2dao2023flashattention2fasterattentionbetter rotary position embeddings (RoPE)su2023roformerenhancedtransformerrotary, RMSNorm huang2019rmsnorm, and SwiGLU activations shazeer2020gluvariantsimprovetransformer. For the Small configuration, we use a hidden size of 512, 6 layers, 8 attention heads, and a feed-forward dimension of 1536. Training is conducted for 6000 steps with a global batch size of 128. We compare AuON, AdamW loshchilov2019decoupledweightdecayregularization, Hybrid-AuON, and MuON under similar training conditions, with learning rates tuned separately: $\eta_{\text{adamw}}=0.003$, $\eta_{\text{auon}}=0.24$. and $\eta_{\text{muon}}=0.01$. rosic2025muon
@@ 4.1 Small Language Modeling
4.2 Comparison of Hybrid-AuON and Muon on FineWebEDU-10B
Setup.
We conduct a controlled paired comparison between Temperature-scaled Hybrid-AuON and Muon on a H100 gpu with a 12-layer GPT-style decoder with hidden size $d=768$ and 6 attention heads(168 Million parameter). Training data are streamed from the 10B-token FineWebEDU kydlicek2025finepdfs corpus using a custom generator that samples contiguous 8k-token sequences while respecting document boundaries via alignment to the GPT-2 end-of-sequence token (50256). Each training step consumes a single 8k-token sequence for both input and target. Validation uses the same sequence length with a fixed budget of 8 sequences (64k tokens), evaluated every 100 steps.
Temperature-scaled variants use
$g_{\text{scaled}}=\gamma\,\tilde{G},\quad\text{with}\quad\gamma=N^{\alpha},$
where $N$ is the tensor size. Larger $\gamma$ amplifies outliers inside $\cosh$, increasing brake strength and $\alpha$ is 0.48.
Vanilla AuON computes
$r=\sqrt{\mathbb{E}[\cosh^{2}(\tilde{G})]}$,
while temperature-scaled variants use
$r=\big(\mathbb{E}[\cosh^{2}(\tilde{G})]\big)^{\beta},\quad\beta=3.5,$
yielding super-linear suppression that engages earlier under spiky gradients.It just change how heard and when the brack engages
Optimization.
Parameters are split into two groups. Scalar parameters, embeddings, and the LM head are optimized with Adam ($\text{lr}=0.008$, $\beta=(0.8,0.95)$, $\varepsilon=10^{-10}$, no weight decay). Hidden matrix parameters in the transformer blocks are optimized with either: (i) Hybrid-AuON, using $\text{lr}=0.048$, momentum =0.95=0.95, weight decay =0.01=0.01, and a 5-step Newton–Schulz orthogonalization applied only to matrix-shaped tensors; or (ii) Muon, using $\text{lr}=0.05$, momentum =0.95=0.95, weight decay =0.01=0.01, with its standard zero-power update. Both share a common schedule $\mathrm{get\_lr}(\mathrm{step})$ consisting of an initial constant phase followed by a linear decay from $1.0$ to $0.1$ over the final $45\%$ of the 5,000 training iterations. A single RNG seed is used so that both runs start from identical initializations and observe the same data order, ensuring a strict paired comparison.
Figure 7: Validation loss on FineWebEDU-10B comparing Hybrid-AuON (blue) and Muon (orange). Both runs share identical random seeds, data streams, and Adam updates on non-matrix parameters. Hybrid-AuON achieves consistently lower validation loss throughout training, particularly in the early and mid stages.
Table 2: Final validation loss after 5,000 training steps. Hybrid-AuON (NS=5) achieves a lower loss than Muon.
Optimizer Configuration Final Val. Loss
Hybrid-AuON Newton–Schulz (5 steps) 4.2882
Muon Zero-power orthogonal update 4.4349
@@ 4.3 Combined Analysis: MNIST and CIFAR-10 CNN
We conduct a combined evaluation of AuON, SGD, and AdamW on two vision benchmarks: a simple multilayer perceptron (MLP) on MNIST and a convolutional neural network (CNN) on CIFAR-10. These experiments, adapted from Wei (2025) wei2025muontutorial, are designed to assess both convergence behavior and runtime efficiency in progressively more challenging tasks.
@@ 5 Conclusion
We introduced AuON, a linear-time optimizer that enforces unit spectral-norm constraints to stabilize training without the computational overhead of full semi-orthogonalization. Across tasks of varying complexity—from nanoGPT on the SmolLM-Corpus to MNIST and CIFAR-10—AuON consistently improved convergence and accuracy over SGD, while matching or surpassing AdamW at nearly identical runtime cost. The hybrid variant (Ns = 5) further outperforms MuON in language modeling, providing a strong and efficient alternative with competitive empirical performance.
Despite these advantages, both AuON and Hybrid-AuON still exhibit attention-logit explosion on very large-scale transformer models, although the effect is less severe than in MuON, as reported in prior work kimiteam2025kimik2openagentic. Techniques such as QK-clipping kimiteam2025kimik2openagentic may help mitigate these issues in future applications.
Overall, our findings show that AuON is a practical and efficient alternative to standard optimizers and a lightweight counterpart to MuON, capable of delivering robust optimization performance without sacrificing computational efficiency in small-scale settings.
@@ Appendix A.1  Convergence Guarantee for AuON
where $N$ is the update dimension.
@@ Appendix A.2  Experimental Validation of Assumptions
Table 6 reports representative values at different training steps. We observe that the training loss decreases monotonically, $\hat{\kappa}$ remains strongly positive (median values between $8.05$ and $13.66$, with 10th-percentiles always above $7$), and $\hat{\sigma}^{2}$ is tightly bounded near $1.0$. This matches the theoretical spectral trust-region guarantee of AuON: updates remain well-aligned while their variance is constrained to a stable unit scale.
@@ 28 U = update / (rms + 1e-8)
33class Auon(torch.optim.Optimizer):